OpenStudy (anonymous):

What is the standard form equation of the line passing through (1, –5) with a slope of –1?

OpenStudy (anonymous):

x+y+4=0

OpenStudy (anonymous):

But The Options Are : x + y = 4 x – y = 4 x – y = –4 x + y = –4

OpenStudy (anonymous):

its x+y=-4

OpenStudy (saifoo.khan):

right.